OpenStudy (anonymous):

What is the equation of the line in slope-intercept form of Y= Use fractions or integers. (-1,-5) and (-3,-2)

OpenStudy (anonymous):

Alright, we have points (-1,5) and (-3,-2) as our given so, plug these in in the general formula of the slope which is:\[m = (y2-y1)/(x2-x1)\]\[m = [(-2)-(-5)]/[(-3)-(-1)]\]\[m = -2+5/-3+1\]\[m = -3/2\] now we have the slope which is m=-3/2, next choose one of the 2 points to use in your equation : y-yo = m(x-xo) choose any point and plug it in :) give it a try now ^_^

OpenStudy (anonymous):

lol sorry, (-1,-5) * ^_^"